# Cross-Covariate Gait Recognition: A Benchmark

### 22 Dec 2023 v1

@@ Introduction
GEINet 2016 4.62 12.7 3.10 9.20
GaitSet 2019 35.3 58.9 25.3 46.7
GaitPart 2020 32.7 55.5 22.6 42.9
GaitGL 2021 35.2 54.1 23.1 39.9
CSTL 2021 11.8 20.1 7.25 13.79
GaitBase 2023 43.8 64.4 31.3 51.3
DeepGaitV2 2023 55.2 75.2 42.5 63.2
GaitGraph 2021 25.2 51.6 15.2 37.2
GaitGraph2 2022 0.27 1.41 0.26 1.4
@@
Parsing Gait (Ours) GaitSet 42.8 67 31.6 54.8
GaitPart 40.9 64.5 29.0 51.5
GaitGL 42.1 61.4 28.4 46.6
CSTL 40.7 61.5 27.9 47.1
GaitBase 56.9 76.0 43.2 63.7
DeepGaitV2 67.2 87.7 52.7 74.7
@@ Acknowledgements
Professor Jin Tang provided financial support for this work. This work was also supported by the National Key Research and Development Program of China under Grant (No.61976144) and the Shenzhen International Research Cooperation Project under Grant (No.GJHZ20220913142611021).

### 05 Jan 2024 v2

@@ Introduction
GEINet 2016 3.10 4.62 9.20 12.7
GaitSet 2019 25.3 35.3 46.7 58.9
GaitPart 2020 22.6 32.7 42.9 55.5
GaitGL 2021 23.1 35.2 39.9 54.1
CSTL 2021 7.25 11.8 13.79 20.1
GaitBase 2023 31.3 43.8 51.3 64.4
DeepGaitV2 2023 42.5 55.2 63.2 75.2
GaitGraph 2021 15.2 25.2 37.2 51.6
GaitGraph2 2022 0.26 0.27 1.4 1.41
@@
Parsing Gait (Ours) GaitSet 31.6 42.8 54.8 67
GaitPart 29.0 40.9 51.5 64.5
GaitGL 28.4 42.1 46.6 61.4
CSTL 27.9 40.7 47.1 61.5
GaitBase 43.2 56.9 63.7 76.0
DeepGaitV2 52.7 67.2 74.7 87.7
@@ Acknowledgements
This work wad supported by the Natural Science Foundation of Hunan Province (No.2023JJ30697), the Changsha Natural Science Foundation (No.kq2208286) and the National Natural Science Foundation of China (No.61502537). This work was also supported in part by the National Key Research, and in part by Development Program of China under Grant (No.61976144) and the Shenzhen International Research Cooperation Project under Grant (No.GJHZ20220913142611021).